# Toward Attention-based TinyML: A Heterogeneous Accelerated Architecture and Automated Deployment Flow

### v1

@@ Abstract
In recent years, Tiny Machine Learning (tinyML) has attracted much attention, bringing compute-intensive Artificial Intelligence (AI) models towards deployment on Microcontroller (MCU) class devices with power envelopes of a few mW. Embedding Deep Neural Networks (DNNs) in small, low-power devices is highly relevant for numerous applications ranging from multi-modal sensing and keyword spotting to anomaly detection and smart wake-up [1]. Compared with cloud-only inference, tinyML offers lower network utilization, higher privacy, and more predictable latency. However, extreme-edge devices typically run with a tightly constrained memory budget, without fully-fledged operating systems and advanced hardware features such as Memory-Management Units (MMUs) and hardware-managed caches.
@@
One of the key research challenges is whether it is possible to build systems that respect the tight hardware and software cost and power constraints of tinyML systems while supporting the rapid advancement of models. A key consideration for addressing this research question is the trade-off between specialization and generality on the computer architecture level. While a breadth of model-specific accelerators has been proposed in recent years [1], making Systems-on-Chip (SoCs) that integrate such accelerators resilient enough to changes in AI models remains an open challenge, especially when memory is tightly constrained. Moreover, automatically and efficiently deploying rapidly evolving DNN models, most notably the now pervasive Attention-based networks, on MCUs enhanced with accelerators, is a major challenge, as product cycles are fast and cannot tolerate the time and cost associated with hand-tuning every model for deployment.
@@
In this paper, we address what we believe to be a fundamental question for the future of tinyML: How can we move from classical perceptive AI and Convolutional Neural Network (CNN) models toward the new leading Attention-based Transformer models? To address this challenge, we leverage a flexible MCU-class architectural template for efficiently integrating specialized hardware accelerators with multi-core clusters over a low-latency Tightly-Coupled Data Memory (TCDM) interconnect. We introduce an extensible deployment flow based on a bottom-up DNN compiler, Deeploy, that enables fast and automated end-to-end deployment. Using the template, we integrate the Integer Transformer Accelerator (ITA) [2] into a RISC-V (RV32) compute cluster based on the Snitch core [3] and prove our hardware-software approach on Attention-based models: as a concrete use case, we showcase the End-to-End (E2E) deployment of the MobileBERT [4] network within a power envelope of (GlobalFoundries fully-depleted silicon-on-insulator technology at ).
@@
We propose a flexible hardware-software architecture template designed to meet the dataflow and compute requirements of emerging Attention-based AI workloads. Our hardware architecture allows the co-integration of multi-core compute clusters with complex hardware accelerators over a high-bandwidth, low-latency TCDM interconnect. At the same time, our co-optimized software template facilitates efficient end-to-end workload mapping. We demonstrate that our hardware-software template enables not only straightforward cooperative execution of workloads between cores, the Direct Memory Access (DMA) engine, and accelerators through shared memory but also starvation-free contention for resources, achieving accelerator utilization of .
As a concrete tinyML use-case, we integrate ITA, a Transformer accelerator, into our hardware template and evaluate its performance through post-layout power analysis, achieving a peak performance of and energy efficiency of .
We demonstrate an E2E deployment flow for MobileBERT on the proposed heterogeneous architecture. With our integration into a flexible DNN compiler, Deeploy, we unlock the potential for collaborative execution between the cluster and the hardware accelerator, which optimizes performance and energy efficiency and prevents resource starvation. By enabling this collaborative execution, we significantly enhance E2E inference energy efficiency by 102 compared to inference without the accelerator, achieving an E2E throughput of up to and energy efficiency of at and .
@@ II-B TinyML Software Deployment Challenges
While several code generation tools for CNNs have been demonstrated [8, 9], most do not generalize to Attention-based models. While CNNs use few branches in their dataflow graphs and therefore do not require sophisticated memory allocation strategies, the highly parallel and branching structure of Attention-based networks requires novel lifetime analysis and tiling strategies to effectively tile and schedule their execution.
@@ III Architecture Template
In this Section, we describe a flexible architecture template, shown in Figure 1, that combines multiple Digital Signal Processing (DSP) optimized RISC-V cores into a compute cluster and facilitates the integration of newly developed hardware accelerators using the HWPE infrastructure. Compared to a single-core system, this enables efficient operation through higher performance and parallelism and enhances adaptability and scalability.The HWPE interface111https://hwpe-doc.readthedocs.io/en/latest/index.html developed for the Parallel Ultra Low Power (PULP) platform [10] facilitates the integration of accelerators with multi-core compute cluster into a shared memory cluster. Our template integrates the area-efficient Snitch cores, occupying each [3]. Snitch is a single-stage, in-order core implementing integer base RV32I, RV32M subset for integer multiply/divide instructions, and standard atomic instruction extension RV32A. Unlike CV32E40P cores used in other PULP-derived clusters222https://docs.openhwgroup.org/projects/cv32e40p-user-manual, Snitch cores are significantly smaller (-56%) and provide a decoupled memory interface, allowing latency-tolerant memory access by pipelining multiple loads and stores.
We couple the cores and accelerators through the shared interleaved L1 TCDM to facilitate energy-efficient data exchange between the compute elements. This is especially crucial for rapidly evolving Attention-based networks as various auxiliary operations need to be computed on the cluster while the majority of the computation is conducted on the accelerator. To reduce banking conflicts and provide the high bandwidth Attention accelerators need, we use 32 banks with each, resulting in a total capacity of . The multi-banked memory makes it unnecessary to attach additional private memory to the accelerator, as data can be accessed by both the accelerator and the cluster’s cores simultaneously. We use a 64-bit TCDM interconnect to enable easy integration of the cluster into a 64-bit host system. The interconnect is implemented as a combinational crossbar, resulting in single-cycle latency in the absence of conflicts with bandwidth towards the L1. Each core has one master port with decoupled request and response path connected to the TCDM interconnect, and the HWPE subsystem features a parametric number of master ports to allow the integration accelerators.
@@ III-A HWPE Subsystem
The HWPE template provides three modules: a controller, one or multiple streamers, and the engine. The controller is the interface between the cores in the cluster and the accelerator. It has a Finite State Machine (FSM) specific to the engine to govern the operation of the accelerator and a memory-mapped register file to keep parameters for the accelerator. The register file can hold a sequence of multiple tasks that can be programmed by each core in the cluster through the controller interface. The streamers act as a special-purpose low-cost DMA to load and store data from the shared TCDM. Finally, the engine contains a hardware accelerator that accepts the streamer’s data and the controller’s configuration.
@@ III-B Neural Network Deployment Framework
To execute Transformer models on the proposed architectural template, we integrate our hardware template in the Deeploy compiler333Deeploy, including our implementation of the proposed hardware template, will be open-sourced on the PULP Platform GitHub (https://github.com/pulp-platform)., which maps neural networks to user-defined, platform-specific C code kernel templates. Deeploy is a DNN compiler that offers architecture-agnostic tinyML optimizations like double-buffering, memory-aware operator tiling, DMA-aware code generation, and fully static offline memory layout generation. These features allow us to accommodate the custom tiling required for operators exclusively present in Transformer networks.
@@
Weights are stored in a double-buffered weight memory to fetch the next set of weights while performing computation with the current set of weights.
Inputs are fetched via streamers and passed through the ITAMax module during step.
While is computed, the ITAMax module operates on the outputs to accumulate the denominator. ITAMax operates in three stages:
Find the local maximum and compare it with the previous maximum stored in the buffer, accumulate the denominator of the Softmax using the current maximum and normalize the previous sum if the maximum is changed.
After the accumulation, the denominator is inverted and saved to the same buffer.
Inputs for step are normalized using the saved maximum and inverted denominator.IV Implementation
As a concrete implementation of our template, we show a platform that couples a cluster with 8+1 RV32IMA Snitch cores with the Integer Transformer Accelerator (ITA) [2]. The ITA accelerator enables the acceleration of 8-bit General Matrix Multiplication (GEMM) and the more complex Multi-Head Attention (MHA) present in Transformer networks.
@@ IV-A Integer Transformer Accelerator (ITA)
With this unique dataflow, ITAMax performs Softmax without additional latency and data fetching from the L1 memory with a low area and power overhead. Since ITA integrates a datapath for Single-Head Attention (SHA), MHA must be calculated sequentially head-by-head. To this end, ITA operates on a single head at a time and computes the partial output projection for each head. The partial outputs of each head need to be summed by the external cluster.
@@ IV-C Physical Implementation
To evaluate our architecture in a tinyML-friendly technology node, we implemented the complete Snitch cluster with the ITA accelerator in GlobalFoundries’ FDX fully-depleted silicon-on-insulator (FD-SOI) technology, targeting an operating frequency of under typical conditions (TT, , ), and in the energy-efficient core voltage configuration (TT, , ). The complete cluster requires (5 MGE) with ITA occupying of the total area. The longest paths of the design are located between the input to the output of the dot product units in the HWPE, within the DMA, and the instruction cache to the data mover core with gate delays of 12, 11, and 11, respectively.
@@ IV-D Neural Network Deployment
To extend Deeploy with our architecture template, including the cluster and ITA, the mapping process of ITA-compatible operators is implemented in a multi-step approach. We start by matching an MHA pattern and fuse it to form a monolithic node in the graph. This node is then split along the head dimension to map the MHA operator head-by-head on ITA. Finally, a head accumulation layer is inserted at the end, which runs on the cluster cores.
@@
To the best of our knowledge, this is the first deployment flow that supports the end-to-end acceleration of Attention-based Transformers at the edge.
@@ V Results
To measure the power consumption and latency of deployed workloads on our design, we perform cycle-accurate post-layout simulation of the entire cluster using Siemens QuestaSim for latency and throughput evaluation at and post-layout gate-level simulations for power measurement under typical conditions (TT, , ). We choose the operating corner to maximize energy efficiency. Our simulation setup accounts for latency and energy costs of memory transfers between the L1 and the system’s background memory via the DMA, programming of the accelerator and cores, and execution of the operators, both on the cluster and ITA. In the following sections, we profile representative microbenchmarks and an end-to-end MobileBERT application. Finally, we compare our results with state-of-the-art MCU-class heterogeneous SoCs for tinyML.
V-A Microbenchmarking Result
We analyze the performance and efficiency of GEMM and the more complex Attention operations and compare the multi-core cluster without any accelerator with the ITA integrated cluster. Calculating GEMMs on ITA offers up to three orders of magnitudes of throughput improvement reaching up to and energy efficiency of corresponding to improvement compared to the cluster without ITA while achieving a peak accelerator utilization of . Running SHA calculation offers an even higher performance improvement of more than 3 orders of magnitudes and a better energy efficiency resulting in and with accelerator utilization. This trend can be attributed to the efficient Softmax implementation in ITA, which does not add latency and thus avoids bottlenecking the overall efficiency.
V-B End-To-End Deployment Results
To benchmark the execution of a complete model, we quantize a MobileBERT network to 8-bit for sentence classification on the SST-2 task of the GLUE benchmark using the QuantLib444https://github.com/pulp-platform/quantlib library. The model has a sequence length , embedding size , projection dimension , parallel Attention heads, layers, and an intermediate size for the feed-forward layer. Due to the extensive simulation time, we measure each layer separately and sum their execution times to extrapolate to the entire network. Table I display the E2E results for MobileBERT for two scenarios: multi-core cluster without the accelerator and multi-core cluster with the ITA accelerator.
| Ours † | Commercial Devices | |||||
@@ V-B End-To-End Deployment Results
| Throughput | [GOp/s] | 0.752 | 154 | 2-7 | 2-45 | 10-60 |
| Energy Efficiency | [GOp/J] | 28.9 | 2960 | 280-400 | 50-560 | 150-650 |
| Power | [mW] | 26.0 | 52.0 | - | - | - |
| Energy per Inference | [mJ] | 164 | 1.60 | - | - | - |
| Inference per Second | [1/s] | 0.16 | 32.5 | - | - | - |
The deployed MobileBERT network requires per inference with sequence length .
@@
In the scenario with a multi-core cluster, using ITA improves throughput at higher energy efficiency.
@@ V-C Comparison with the State-of-the-art
The Syntiant NDP120555https://www.syntiant.com/hardware MCU implemented in UMC ULP technology uses the Syntiant Core 2 tensor processor coupled with an Arm Cortex M0 processor and a HiFi-3 DSP. It achieves up to at in MLPerf Tiny Inference on MobileNetV1 [8]. We also compare with the Ensemble E3 AI MCU from Alif Semiconductor666https://alifsemi.com/ which couples Ethos-U55 Machine Learning (ML) processors with ARM Cortex M55 processors. Depending on the network it achieves up to at . Compared to both devices, we achieve at least more throughput with a higher energy efficiency.

### v2

@@ Abstract
In recent years, Tiny Machine Learning (tinyML) has attracted much attention, bringing compute-intensive Artificial Intelligence (AI) models towards deployment on Microcontroller (MCU) class devices with power envelopes of a few milliwatts. Embedding Deep Neural Networks (DNNs) in small, low-power devices is highly relevant for numerous applications ranging from multi-modal sensing and keyword spotting to anomaly detection and smart wake-up [1]. Compared with cloud-only inference, tinyML offers lower network utilization, higher privacy, and more predictable latency. However, extreme-edge devices typically run with a tightly constrained memory budget, without fully-fledged operating systems and advanced hardware features such as Memory-Management Units (MMUs) and fully automated cache hierarchies.
@@
One of the key research challenges is whether it is possible to build systems that respect the tight hardware and software cost and power constraints of tinyML systems while supporting the rapid advancement of models. A key consideration for addressing this research question is the trade-off between specialization and generality on the computer architecture level. Although numerous model-specific accelerators have been proposed in recent years [1], designing Systems-on-Chip (SoCs) that can integrate these accelerators while remaining adaptable to evolving AI models remains an open challenge, particularly under tight memory constraints. Moreover, automatically and efficiently deploying rapidly evolving DNN models, especially the increasingly popular Attention-based networks, on accelerator-enhanced MCUs remains a significant challenge. Additionally, fast product cycles make it difficult to accommodate the time and cost associated with hand-tuning each model for deployment.
@@
In this paper, we address what we believe to be a fundamental question for the future of tinyML: How can we move from classical perceptive AI and Convolutional Neural Network (CNN) models toward leading Attention-based Transformer models? Unlike in CNNs, complex dataflow operations like Softmax in Transformers can lead to high latency despite their low arithmetic complexity. While General Matrix Multiplication (GEMM) accelerators handle most computations in Transformer networks efficiently, the remaining operations can become a bottleneck. To address this challenge, we leverage a flexible MCU-class architectural template for efficiently integrating specialized hardware accelerators with multi-core clusters over a low-latency Tightly-Coupled Data Memory (TCDM) interconnect. At its core, we use a RISC-V (RV32) compute cluster based on the latency-tolerant Snitch core [2]. To the best of our knowledge, this is the first heterogeneous Snitch-based cluster integrating Hardware Processing Engines (HWPEs) 111https://hwpe-doc.readthedocs.io/en/latest/index.html, advancing beyond previous configurations which focused on instruction extension units tightly coupled to the pipeline. We introduce an extensible deployment flow based on a bottom-up DNN compiler, Deeploy, that enables fast and automated End-to-End (E2E) deployment. Using this template, we integrate an extended version of the Integer Transformer Accelerator (ITA) [3] and prove our hardware-software co-design flow on Attention-based models. As a concrete use case, we showcase the E2E deployment of MobileBERT [4], DINOv2 [5], and Whisper’s encoder [6] within a power envelope of (GlobalFoundries fully-depleted silicon-on-insulator technology at ).
@@
We propose a novel, flexible hardware-software architecture template designed to meet the dataflow and compute requirements of emerging Attention-based AI workloads. Our hardware architecture allows the co-integration of a multi-core latency-tolerant Snitch compute cluster with complex hardware accelerators over a high-bandwidth, low-latency TCDM interconnect. At the same time, our co-optimized software template facilitates efficient E2E workload mapping. We demonstrate that our hardware-software template enables starvation-free contention for resources in the shared memory with its tunable interconnect bandwidth and the Direct Memory Access (DMA) engine. As a result, we achieve accelerator utilization of up to .
As a concrete use-case, we integrate ITA, a Transformer accelerator tuned for the specific dataflow of the Attention calculation, into our hardware-software template and extend Deeploy222https://github.com/pulp-platform/Deeploy with an accelerator model to enable automated mapping, scheduling, tiling, and code generation. We evaluate the performance through post-layout power analysis, achieving a peak performance of and energy efficiency of up to . The integration incurs only a decrease in utilization compared to the standalone accelerator, demonstrating the low overhead of our template.
We showcase the capability of our hardware-software template to support a range of Attention-based tinyML models, including MobileBERT, DINOv2, and Whisper’s encoder. Our flow unlocks the potential for collaborative execution between the cluster and the hardware accelerator, which optimizes performance and energy efficiency and prevents resource starvation. By enabling this collaborative execution, we significantly enhance E2E inference energy efficiency by 102 compared to inference without the accelerator, achieving an E2E throughput of up to and energy efficiency of .
@@ II-B TinyML Software Deployment Challenges
While several code generation tools for CNNs have been demonstrated [1], most do not generalize to Attention-based models. While CNNs use few branches in their dataflow graphs and therefore do not require sophisticated memory allocation strategies, the highly parallel and branching structure of Attention-based networks requires novel lifetime analysis and tiling strategies to effectively tile and schedule their execution.
@@ III Architecture Template
In this Section, we describe a flexible architecture template, shown in Figure 1, that combines multiple Digital Signal Processing (DSP) optimized RISC-V cores into a compute cluster and facilitates the integration of newly developed hardware accelerators using the HWPE infrastructure and automated deployment. Compared to a single-core system, this enables efficient operation through higher performance and parallelism and enhances adaptability and scalability. The HWPE interface developed for the Parallel Ultra Low Power (PULP) platform facilitates the integration of accelerators with multi-core compute cluster into a shared memory cluster. Our template integrates the area-efficient Snitch cores, occupying each [2]. Snitch is a single-stage, in-order core implementing integer base RV32I, RV32M subset for integer multiply/divide instructions, and standard atomic instruction extension RV32A. Unlike CV32E40P cores used in other PULP-derived clusters333https://docs.openhwgroup.org/projects/cv32e40p-user-manual, Snitch cores are significantly smaller (-56%) and provide a decoupled memory interface, allowing latency-tolerant memory access by pipelining multiple loads and stores.
We couple the cores and accelerators through the shared interleaved L1 TCDM to facilitate energy-efficient data exchange between the compute elements. This is especially crucial for rapidly evolving Attention-based networks as various auxiliary operations need to be computed on the cluster while the majority of the computation is conducted on the accelerator. To reduce banking conflicts and provide the high bandwidth Attention accelerators need, we use 32 banks with each, resulting in a total capacity of . The multi-banked memory makes it unnecessary to attach additional private memory to the accelerator, as data can be accessed by both the accelerator and the cluster’s cores simultaneously. We use a 64-bit TCDM interconnect, which is implemented as a combinatorial crossbar, resulting in single-cycle latency in the absence of conflicts with bandwidth towards the L1. Each core has one master port with decoupled request and response path connected to the TCDM interconnect, and the HWPE subsystem features a parametric number of master ports to allow the integration accelerators.
@@ III-A HWPE Subsystem
The HWPE template provides three modules: a controller, one or multiple streamers, and the engine. The controller is the interface between the cores in the cluster and the accelerator. It has a Finite State Machine (FSM) specific to the engine to govern the operation of the accelerator and a memory-mapped register file to keep parameters for the accelerator. The register file can hold a sequence of multiple tasks that can be programmed by any core in the cluster through the controller interface over the narrow AXI interconnect. A task represents a set of configuration values used by the accelerator. The streamers act as a special-purpose low-cost DMA to load and store data from the shared TCDM. Finally, the engine contains a hardware accelerator that accepts the streamer’s data and the controller’s configuration.
@@ III-B Neural Network Deployment Framework
To execute Transformer models on the proposed architectural template, we integrate our hardware template in the Deeploy compiler [9], which maps neural networks to user-defined, platform-specific C code kernel templates. Deeploy is a DNN compiler that offers architecture-agnostic tinyML optimizations like double-buffering, memory-aware operator tiling, DMA-aware code generation, and fully static offline memory layout generation. These features allow us to accommodate the custom tiling required for operators exclusively present in Transformer networks.
@@
IV Implementation
As a concrete implementation of our template, we show a platform that couples a cluster with 8+1 RV32IMA Snitch cores with the Integer Transformer Accelerator (ITA) [3]. The ITA accelerator enables the acceleration of 8-bit GEMM and the more complex multi-head Attention (MHA) present in Transformer networks. ITA used in this work is an extended version of the accelerator presented in [3], featuring additional functionality through the inclusion of a partial sum buffer and an activation unit supporting ReLU and GeLU. Furthermore, it is wrapped with HWPE components.
@@ IV-A Integer Transformer Accelerator (ITA)
With this unique dataflow, ITAMax performs Softmax without additional latency and data fetching from the L1 memory with a low area and power overhead. Since ITA integrates a datapath for single-head Attention, MHA must be calculated sequentially head-by-head. Therefore, ITA operates on a single head at a time and computes the partial output projection for each head. The partial outputs of each head need to be summed by the external cluster.
@@ IV-C Physical Implementation
To evaluate our architecture in a tinyML-friendly technology node, we implemented the complete Snitch cluster with an extended version of the ITA accelerator in GlobalFoundries’ FDX fully-depleted silicon-on-insulator (FD-SOI) technology, targeting an operating frequency of under typical conditions (TT, , ), and in the energy-efficient core voltage configuration (TT, , ). The extended design includes a partial sum buffer, an activation unit, and the HWPE components. The complete cluster requires (5 MGE) with the HWPE subsystem occupying of the total area. The longest paths of the design are located between the input to the output of the dot product units in the HWPE, within the DMA, and the instruction cache to the data mover core with gate delays of 12, 11, and 11, respectively.
@@ IV-D Neural Network Deployment
To extend Deeploy with our architecture template, including the cluster and ITA, the mapping process of ITA-compatible operators is implemented in a multi-step approach. Deeploy starts by matching an MHA pattern and fuses it to form a monolithic node in the graph. This node is then split along the head dimension to map the MHA operator head-by-head on ITA. Finally, a head accumulation layer is inserted at the end, which runs on the cluster cores.
@@
To the best of our knowledge, this is the first deployment flow that supports the E2E acceleration of Attention-based Transformers at the edge.
@@ V Results
Ours | Commercial Devices | |||||
@@
Throughput | [GOp/s] | 0.74 | 56-154 | 2-7 | 2-45 | 10-60 |
Energy Efficiency | [GOp/J] | 28.9 | 1600-2960 | 280-400 | 50-560 | 150-650 |
Power | [mW] | 26.0 | 35.2-52.0 | - | - | - |
MobileBERT a | DINOv2-Small b | Whisper-Tiny Encoder c | |||||
Metric | Unit | Multi-Core | Multi-Core + ITA | Multi-Core | Multi-Core + ITA | Multi-Core | Multi-Core + ITA |
Energy per Inference | [mJ/Inf] | 164 | 1.60 | 407 | 7.31 | 340 | 5.55 |
Inference per Second | [Inf/s] | 0.16 | 32.5 | 0.06 | 4.83 | 0.08 | 6.52 |
@@
a
per inference with sequence length
b
per inference with sequence length
c
per inference with sequence length
To measure the power consumption and latency of deployed workloads on our design, we perform cycle-accurate post-layout simulation of the entire cluster using Siemens QuestaSim for latency and throughput evaluation at and post-layout gate-level simulations for power measurement under typical conditions (TT, , ). We choose the operating corner to maximize energy efficiency. Our simulation setup accounts for latency and energy costs of memory transfers between the L1 and the system’s background memory via the DMA, programming of the accelerator and cores, and execution of the operators, both on the cluster and ITA. In the following sections, we profile representative microbenchmarks and the execution of three different Transformer networks. Finally, we compare our results with state-of-the-art MCU-class heterogeneous SoCs for tinyML.
V-A Microbenchmarking Result
We analyze the performance and efficiency of GEMM and the more complex Attention operations and compare the multi-core cluster without any accelerator with the ITA integrated cluster. Our heterogenous cluster achieves a throughput of and energy efficiency of in GEMM computation, corresponding to and improvement respectively compared to the cluster without ITA with a peak accelerator utilization of . Running single-head Attention operation offers an even higher performance improvement of more than 3 orders of magnitudes and a better energy efficiency resulting in and with accelerator utilization. The standalone accelerator achieves a slightly higher utilization of , with the integration into the template incurring only a small decrease of . This demonstrates that the template has minimal impact on the accelerator utilization. This trend can be attributed to the efficient Softmax implementation in ITA, which does not add latency and thus avoids bottlenecking the overall efficiency.
V-B End-To-End Deployment Results
To benchmark the execution of a complete model, we quantize MobileBERT444, , , , , (Sequence Length, Embedding Size, Projection Dimension, Attention Heads, Layers, Feed-Forward), DINOv2555, , , , , and Whisper’s666, , , , , encoder using the QuantLib777https://github.com/pulp-platform/quantlib library to perform 8-bit full integer inference. Due to the extensive simulation time, we measure each layer separately and sum their execution times to extrapolate to the entire network. Table I display the E2E results for two scenarios: multi-core cluster without the accelerator and multi-core cluster with the ITA accelerator.
In the scenario with a multi-core cluster, using ITA improves throughput up to at higher energy efficiency.
@@ V-C Comparison with the State-of-the-art
The Syntiant NDP120888https://www.syntiant.com/hardware MCU implemented in UMC ULP technology uses the Syntiant Core 2 tensor processor coupled with an Arm Cortex M0 processor and a HiFi-3 DSP. It achieves up to at in MLPerf Tiny Inference on MobileNetV1 [11]. We also compare with the Ensemble E3 AI MCU from Alif Semiconductor999https://alifsemi.com/ which couples Ethos-U55 Machine Learning (ML) processors with ARM Cortex M55 processors. Depending on the network it achieves up to at . Compared to both devices, we achieve at least more throughput with a higher energy efficiency.